OpenStudy (anonymous):

HELP NEEDED -3x + y = -3 x + y = 1 b. x + y = 1 x- 3y = -3 c. 3x + 2y = 6 -3x + 4y = 12 d. x + y = 5 x - 3y = -3 How do you solve these to get the ordered pair so I can graph them?

OpenStudy (anonymous):

@hashbasl

OpenStudy (anonymous):

You mean to get them into standard form (mx + b)?

OpenStudy (anonymous):

no, i want to get them into an ordered pair. like (3,5) I honestly have no idea what i am doing.

OpenStudy (anonymous):

Oh! You want to find the solution :)

OpenStudy (anonymous):

YESSS!!!!! that was the word that i was looking for. i have graphs too. the graphs are the only thing i can really read. lol

OpenStudy (anonymous):

The solution to a system of equations is looks like this on the graph 

OpenStudy (anonymous):

Its where they intersect

OpenStudy (anonymous):

Do you want to know how to solve this algebraically?

OpenStudy (anonymous):

gotcha...

OpenStudy (anonymous):

I don't know... 8th grade math, suck at it.

OpenStudy (anonymous):

Well you might need to know how to do it. So lets go through it together.

OpenStudy (anonymous):

OK.

OpenStudy (anonymous):

Say we had x+y = 2 and 2x-y = 5

OpenStudy (anonymous):

We cant really solve one of them by themselves because there is two variables right?

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

So this is what we do, we take one of the equations x + y = 2 x + y - y =2 - y x= 2 -y And we solve for one of the variables! Now we know what x equals. (But we arent done yet)

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):

So our equations were x+y = 2 and 2x-y = 5 right?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Now that we know what x equals we can do something known as substitution.

OpenStudy (anonymous):

what i really need to do is this.

OpenStudy (anonymous):

So if I was to do it to 2x-y = 5 it would look like this 2x - y = 5 and since x = 2 - y we can ... 2(2 - y) - y = 5

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

You see how there is only one variable in the equation now? Now we can solve for the y coordinate of the solution.

OpenStudy (anonymous):

In (x,y)

OpenStudy (anonymous):

this is what i am trying to do

OpenStudy (anonymous):

that is what i am trying to figure out.

OpenStudy (anonymous):

Are you allowed to graph?

OpenStudy (anonymous):

that is what the actual teacher prefers.

OpenStudy (anonymous):

Oh! Wow, that is so much faster.

OpenStudy (anonymous):

I don't know how to manipulate microsoft office.

OpenStudy (anonymous):

Here is a link to a graphing calculator: https://www.desmos.com/calculator

OpenStudy (anonymous):

Type in the equations and match it to the right one

OpenStudy (anonymous):

HOLY CRAP!!!!!! why am i not allowed to know that? if only i knew how to give people medals.

OpenStudy (anonymous):

Click on the blue best response button next to the medal

OpenStudy (anonymous):

Thank you.